An airport is looking for some devices to support the luggage transportation services. There are 4 alternatives available (D1, D2, D3 and D4). To assist the evaluations, their purchase costs, service lives, possible salvages at the end of lives have been collected in Table Q2. In addition, the monthly variable costs of using these devices to perform the luggage transportation services in particular, and the numbers of corresponding devices needed are also provided in Table Q2. Given that the internal rate of return is set to 0.8% per month in that case.

  1. (a) What are the Present Values of the end of life salvages (per unit) from these 4 devices? (4 marks)

  2. (b) Determine the Capital Recovery Factors for the calculations of the monthly equivalent fixed costs (per unit) of the 4 devices. (4 marks)

  3. (c) What are the fixed costs (per unit) in a monthly basis of the 4 devices? (4 marks)

  4. (d) Prioritize the preference of choosing these 4 devices based on the cost minimization purpose and order them in the format of A > B > ..., where A is better than B and so on.

    (8 marks)

  5. (e) What is the total monthly cost of the most preferred choice for providing the luggage

transportation services?

(5 marks)

Rubrics and Marking:

Table Q2

Device

D1

D2

D3

D4

Purchase Cost ($/unit)

38000

18000

35000

12000

Service Life (years)

5

7

8

4

End of Life Salvage ($/unit)

2500

2000

3200

3000

No. of Devices Required (units)

1

2

1

3

Variable Cost Required ($/month/unit)

450

220

380

150
Answer #1
Device D1 D2 D3 D4
Purchase Cost ($/unit) 38000 18000 35000 12000
Service Life (years) 5 7 8 4
                         in months(yrs.*12) 60 84 96 48
End of Life Salvage ($/unit) 2500 2000 3200 3000
No. of Devices Required (units) 1 2 1 3
Variable Cost Required ($/month/unit) 450 220 380 150
Total variable cost(per unit*no.of units) 450 440 380 450
PV of annuity of variable costs
Device D1 D2 D3 D4
A.Purchase Cost ($/unit) 38000 18000 35000 12000
1.Service Life (years) 5 7 8 4
   2.                      in months(yrs.*12) 60 84 96 48
3.End of Life Salvage ($/unit) 2500 2000 3200 3000
(a) Present Values of the end of life salvages (per unit) (Row 3/1.008^Row 2) 2500/1.008^60= 2000/1.008^84= 3200/1.008^96= 3000/1.008^48=
1549.92 1024.11 1489.15 2046.52
4.No. of Devices Required (units) 1 2 1 3
5.Variable Cost Required ($/month/unit) 450 220 380 150
6.Total variable cost for the device 450 440 380 450
7.Monthly Capital Recovery Factors (Ans.b) (1-1.008^-60)/0.008= (1-1.008^-84)/0.008= (1-1.008^-96)/0.008= (1-1.008^-48)/0.008=
47.50421418 60.99331783 66.83001785 39.72838826
8.PV of variable costs for the device(6*7)(Ans.c) 21376.89638 26837.05985 25395.40678 17877.77472
9.Net PV of costs(A-a+8) 57826.98 43812.95 58906.26 27831.26
10.Equivalent monthly cost(NPV/Factor)(9/7) 1217.30 718.32 881.43 700.54
d..so, from the above
from cost minimisation point of view,
D4 > D2> D3 > D1
(e) Total monthly cost of the most preferred choice for providing the luggage = $ 700.54 (Uniform Equivalent Monthly Cost for the device D4)